1. two sides of triangle are the same length.the third side is 6 in. less than the sum of the equal sides.if the perimeter of the triangle is 46 in. find the three sides.
2. the length of a rectangular garden is 12 ft less than four times the width.if the perimeter of the garden is 256 ft. find the dimension of the garden.

Answer by addingup(3677)

Let's call the two equal sides 2s. And the third side is 2s-6:
2s+2s-6 = 46 add the s on the left and add 6 to both sides:
4s= 52 divide both sides by 4
s= 13 this is the length of each of the same length sides. And the 3rd side is:
2(13)-6= 20
Check: 2(13)+2(13)-6= 26+26-6= 46 Correct.
--------------------------------
L is 12 less than 4W. And the perimeter is:
2L+2W= 256 And we just said that L= 4W-12, so let's substitute:
2(4W-12)+2W= 256
8W- 24+ 2W= 256 Now add 24 on both sides and add the W onthe left:
10W = 280 Divide both sides by 10
W= 28 is the width and the Length is:
28*4-12= 100
Check: The perimeter is 2L+2W = 2(100)+2(28)= 200+58= 258 Correct.